# Time-Aware and Transition-Semantic Graph Neural Networks for Interpretable Predictive Business Process Monitoring  [^1] The full codebase is available at https://github.com/skyocean/TemporalAwareGNNs-NextEvent

### v1

@@ 8 Declaration of generative AI and AI-assisted technologies in the writing process
(i) Dataset: Helpdeask; Model: GAT-TD
(j) Dataset: Helpdesk; Model GAT-TDTE
Figure Appendix C.1: Node Importance Heatmap by Sequence Length and Position
@@
(i) Dataset: Helpdeask; Model: GAT-TD
(j) Dataset: Helpdesk; Model GAT-TDTE
(a) Dataset: BPI12; Model: GAT-TD
(b) Dataset: BPI12; Model GAT-TDTE;
(c) Dataset: BPI12W; Model: GAT-TD
(d) Dataset: BPI12W; Model GAT-TDTE
(e) Dataset: BPI13i; Model: GAT-TD
(f) Dataset: BPI13i; Model GAT-TDTE
(g) Dataset: BPI13c; Model: GAT-TD
(h) Dataset: BPI13c; Model GAT-TDTE
(i) Dataset: Helpdeask; Model: GAT-TD
(j) Dataset: Helpdesk; Model GAT-TDTE
Figure Appendix C.4: Top Node Dominance by Rank and length Range

### v2

@@ 8 Declaration of generative AI and AI-assisted technologies in the writing process
Figure Appendix C.1: Node Importance Heatmap by Sequence Length and Position (part 1)
(a) Dataset: Helpdeask; Model: GAT-TD
(b) Dataset: Helpdesk; Model GAT-TDTE
(c) Dataset: BPI20; Model: GAT-TD
(d) Dataset: BPI20; Model GAT-TDTE
Figure Appendix C.2: Node Importance Heatmap by Sequence Length and Position (part 2)
(a) Dataset: BPI12; Model: GAT-TD
(b) Dataset: BPI12; Model GAT-TDTE;
(c) Dataset: BPI12W; Model: GAT-TD
(d) Dataset: BPI12W; Model GAT-TDTE
Figure Appendix C.3: Critical Windows Identification across Length Ranges (part 1)
(a) Dataset: BPI13i; Model: GAT-TD
(b) Dataset: BPI13i; Model GAT-TDTE
(c) Dataset: BPI13c; Model: GAT-TD
(d) Dataset: BPI13c; Model GAT-TDTE
(e) Dataset: Helpdeask; Model: GAT-TD
(f) Dataset: Helpdesk; Model GAT-TDTE
(g) Dataset: BPI20; Model: GAT-TD
(h) Dataset: BPI20; Model GAT-TDTE
Figure Appendix C.4: Critical Windows Identification across Length Ranges (part 2)
@@
Figure Appendix C.6: Top Node Dominance by Rank and length Range (part 1)
(a) Dataset: Helpdeask; Model: GAT-TD
(b) Dataset: Helpdesk; Model GAT-TDTE
(c) Dataset: BPI20; Model: GAT-TD
(d) Dataset: BPI20; Model GAT-TDTE
Figure Appendix C.7: Top Node Dominance by Rank and length Range (part 2)
@@
(a) Dynamic Temporal Windows with Attention Decomposition
Figure Appendix C.13: Temporal Attention Decomposition for a Representative Sequence (9 events) from BPI20 Using GAT-TD and GAT-TDTE Model